Monday, January 31, 2011

More than just raft races

To all those wondering how there can possibly be time for fun stuff like raft races when there's a big telescope to get going - fear not!  The team's slaving away, as always, & much progress is being made...


Shortly before Xmas, Phil was up at SALT & got to try out his fantastic new 10-24 mm lens while Jonathan was up on the cherry-picker to remove the central segment for re-coating.


This was historic stuff as it's the first time that Segment 1, the critical mirror upon which the procedure for aligning the primary hinges, was re-aluminised. 


With that, Hitesh can at last sleep peacefully again, knowing that the nasty little circle hand-drawn on the centre of the segment 3.5 years ago is Gone!


In the time since, much more rain has produced a greener Sutherland than most of us have ever seen before.


& if that wasn't exciting enough, there was the man-hunt for the French couple at large in the Sutherland area after an altercation that left one policeman dead & another injured :(  So much for peaceful, boring old Sutherland!


That story ended badly for the fugitives a few days later.  For the record, we'd like to confirm that the controversially named "gun turret" on RSS (which houses the grating rotation mechanism) played no part in the shoot-out - it really was in the spectrometer room the whole time...


In fact - were we not still waiting on SALTICAM (which too is progressing well down in Cape Town), RSS might soon have been allowed back up on the tracker after its excellent behaviour during exhaustive repetitive testing of the various mechanisms.


The other critical instrument that needs to go into the payload before RSS can go back on top is the Atmospheric Dispersion Compensator (ADC).  It will be inserted into the non-rotating structure (just above the SAC) using a custom-built installation jig that seems destined to spend at least part of its time closely associated with potjiekos events.


All the while, primary mirror segments continue to be cleaned & re-aluminised in the IQ Team's old haunt.


One of the ongoing night-time challenges is to sort out stray light issues due to the auto-collimator, as well as other spurious light sources like LEDs on the payload & within the telescope enclosure.  A large baffle that surrounds the SAC has been installed & aligned, the auto-collimator had its baffle (affectionately known as "the mother of all tubes") replaced & all the LEDs will ultimately need to be hunted down & blacked out.

Nightlog 2011-01-29

SA: Tim
SO: Siphelo
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


another smooth night with good, photometric weather and decent seeing for the most part. got lots of thruput data and a couple data sets with 0.03 sec exposures to do some turbulence analysis. had some tracker issues with motor H5 overheating which slowed us down a little bit, but we were able to work through it. the mass-dimm even worked flawlessly all night. MASS free seeing when i shut down at dawn was 0.12"....

-------------------------------------------

Nightlog 2011-01-30

SA: Tim
SO: Siphelo
Others: Charl

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


got a little bit more thruput data and one round of short exposure burst mirror data. struggled most of the rest of the night trying to get the mirror to behave. several segments were way out of whack from all of our bursting and anti-bursting. one was focusing almost an arcminute away from where it was supposed to. when we finally got all of that sorted out, the humidity came in and closed us up.

-------------------------------------------

Nightlog 2011-01-28

SA: Tim
SO: Siphelo
Others: Charl

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


night didn't look promising at the start, but turned out to be beautiful. warm, clear, light winds, and half-decent seeing for a stretch. got some good practice setting up burst mirror modes and recovering from them as well as recovering from other mirror issues. thanks to charl for coming up to help and to hitesh for putting up with us calling. got some data in burst mode with short exposure times to try and measure seeing/turbulence. got some thruput data at the end with the VI-apogee and its V filter

-------------------------------------------

Nightlog 2011-01-24

SA: Paul
SO: Patrick
Others: Siphelo, Zolisa, Darragh, Lisa, Charl, Hitesh

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- The SAC baffle was reinstalled today. From the out of focus images we seemed
to see multiple spider components which we thought was due to the SAC baffle.
Took the SAC baffle off, but the multiple component remained. This cannot be
explained. The SAC baffled seemed well enough aligned in the way it was
installed. Multi component spider needs thinking about. The SAC baffle will be
re-installed tomorrow during the day.

- timdimm ran for a while but it was plagued by tracking issues.

- we continued the training and focused on getting the mirror alignment times
down to around 15min.

- some pointing offset observations were done and the results from that looks
good.

------------------------------------------------------

Nightlog 2011-01-27

SA: Tim
SO: Siphelo
Others: Charl, Zolisa, Martin

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


smooth, productive night with good weather and mediocre seeing. spent the first half of the night doing stray light tests with a closed dome. spent the second half doing another track following out-of-focus images and got some open-loop tracking data at the end. the mass-dimm ran flawlessly until almost 0400. it somehow quit tracking and lost the stars. updated the commissioning wiki with some of the results from tonight.

-------------------------------------------

Nightlog 2011-01-26

SA: Tim
SO: Siphelo
Others: Charl, Zolisa

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


used BCAM to take some time-lapse data of out-of-focus images to watch the evolution of the mirror alignment over time. the stars used for this came up at the very edge of the BCAM field when slewed to so there's something screwy with the pointing. will need to follow up on that.

made movies out of the time-lapse data and wrote up a blog post on them at http://saltastro.blogspot.com/2011/01/salt-pupil-time-lapse.html.

got closed by humidity right after midnight and ridge cloud by 3 AM.

-----------------------------------------

Nightlog 2011-01-25

SA: Paul
SO: Patrick
Others: Siphelo, Zolisa, Darragh, Lisa, Charl, Hitesh, Tim

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Did some stray light observations see Darragh's email : Autocollimator stray
light for the results obtained.

- timdimm and the MASS was operational tonight. Garith sorted out the
networking issues this afternoon and Tim fixed the loose instrument mount on
the back of the Meade.

- The end of the night was spend on some more pointing offset observations.

------------------------------------------------------

Nightlog 2011-01-23

SA: Paul
SO: Patrick
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- CCAS not operational.

------------------------------------------------------

Thursday, January 27, 2011

SALT Pupil Time-Lapse

While taking some out-of-focus images to check on some vignetting issues we thought it might be interesting to take a sequence of such images to track how the segments behave after an alignment is done. The first such sequence was done while observing the star HR 806. A total of 100 images were taken with 0.1 exposure time each and a 30 second delay between each one. Here is the resulting movie:

video

This particular track didn't allow us to see some parts of the primary so we tried again on a star (HR 1960) that allowed us to track through the center of the primary. We set out to take 200 images (0.5 sec exposure; 30 seconds apart), but the sequence got cut short to 175 when the humidity spiked and forced us to close. Here are the results:


video

Sunday, January 23, 2011

Tech Ops Raft Race Day

After a long night of furious construction in the workshop, the 6 teams & their rafts - in various stages of construction - were relocated to a nearby farm dam on Saturday morning.  The judge's bank account was open for business...


& a flock of spectators rolled up to follow the action.


Some teams seemed hopelessly un-prepared, arriving with little more than their raw materials.


Others had made a start but still had a long way to go.


& some had clearly been extremely busy over the past few days.


"Basic tools" like a welder were available out at the dam...


Even the most mature design still needed substantial development through the course of the morning.


While others hauled out their secret weapon - an outboard motor alledged to run on ethanol...  Fortunately the judge was having none of that & the offending accessory was disallowed.


Slowly but surely - the various craft took shape.


Some were clearly subject to far more creativity than most: enter Johan, Nicolaas, Cobus & Co's Pirate Astronomers team!


A team bringing credibility to astronomers at last :)  This really made an impact on Judge Darragh.


After a few spectacular wipe-outs, the team of Chris, Eben & Hitesh got it together & showed everyone how it's done.


Zulu, Amanda & Jonathan's raft put up more of a fight - throwing them off a few times before making its way to a quiet corner of the dam, from where it later watched the race.


Preliminary sea-trials for the Pirates were fairly harrowing, but they weathered the storm & made the necessary mods to ensure future stability.


Stripped of their first-choice propulsion mechanism, the team of Vic, Peter & Martin had to lighten their raft & start scrounging for paddles.  All the while - Zulu & Amanda wrestled bravely with their recalcitrant beast.


Francois, Grant & Garith's masterpiece was rapidly taking shape.


But progress was relatively slow for Charl, Paul & Lisa's monstrosity with the crank from hell & the super-sized rudder.  


Elegance did not feature prominently in this design, but at least it did eventually come together & let the record reflect that it did in fact float...


In the meantime, the ergonomic, aesthetically pleasing craft-next-door sat patiently anchored, waiting for the show to begin - subliminally taunting the rest of the field.


Once all had survived their sea-trials, it was time to go!


The fleet first had to cross the dam to pick up tokens on the far side before heading over to the finish near the dam "wall".


The field quickly opened up & the combination of Hitesh's peddling & Eben's paddling sped them to the finish. 


Chris was conspicuous in his absence during the business part of the race, but did later enter the water to participate in their lap of honour.  Tsk Tsk!


Francois, Garith & Grant were next in, their sail taking full advantage of the wind.


While all this was going on - an epic drama was unfolding remarkably close to the start...  It's not yet clear whether a broken cable tie, or a powerful vortex generated by the propeller attached to Hitesh's bicycle, led to the spectacular disintegration of the cumbersome paddle steamer.


None the less, the brave crew heroically coaxed their mangled craft (& battered bodies) to the finish - albeit bypassing the check-point on the far side of the dam.  Mercifully, Amanda was on hand to administer first aid in the form of Windhoeks for all - a sign of true compassion!


The team responsible for numerous acts of sabotage throughout the day somehow got hold of a token at the check-point & then sent Martin racing to the finish - on foot (!) - in order to claim 3rd place.  A glaring omission in the rules prevented this being declared illegal & thus the deserving Pirate Astronomers were robbed of a podium finish.


Luckily, by this stage - any sense of competitive urgency had long since dissipated.


Sadly for our judge du jour, cellphone reception was patchy & did not permit the banking-by-phone that he'd been hoping for.


The winning craft, parked back at the start, exuded an air of cool smugness for the rest of the afternoon.


For the spectators, it was time for a cruise with skipper Eben.


Safely back on terra firma, Charl took it upon himself to administer the obligatory "mampoer" shots to all in need of antiseptic, anaesthetic, solace & rehydration.



The braai was handled by Kobus & helpers - thank you guys!  After expending all that physical, mental & emotional energy, everyone was ravenously hungry by mid-afternoon.



Next it was time for the prize-giving, with under-sized T-shirts & vouchers for "dinners with Vic" being awarded.


In addition to the 1st, 2nd & 3rd place prizes - the Pirate Astronomers claimed the "most innovative raft" award & Amanda, Jono & Zulu got the "best team spirit" prize for not getting too bummed when their raft threw them off & ran away.


However, as is so often the case - the unsung hero of the day was undoubtedly Charl: the only person to be dealt the crippling blow of having 2 astronomers put into his team.  He also stoically guided his sinking craft to the finish, man-hauled the wreckage back to the start & then helped to provide food & strong alcohol for all - Cheers Boet!


What a fantastic day!  Thanks especially to Chris, Adelaide, Grant & their many helpers for making it happen, to the farmers that offered us the use of their dam & to everyone for their superb contributions to this brilliant event :)

Nightlog 2011-01-22

SA: Paul
SO: Patrick
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- It was cloudy at the start of the evening and cleared later accompanied by
high humidity.

- The CCAS instrument has been taken apart for testing.

------------------------------------------------------

Saturday, January 22, 2011

Nightlog 2011-01-21

SA: Paul
SO: Patrick
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Its been raining the whole evening, did not open.

------------------------------------------------------

Thursday, January 20, 2011

Nightlog 2011-01-19

SA: Paul
SO: Patrick
Others: Zolisa, Siphelo, Garith, Charl

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Did not open the dimms due to the high wind.

- Charl gave the SO's some more tips and tricks on getting the mirror alignment
times down and discussed some of the pitfalls. We had the perfect conditions to
do just that, a strong wind and a large temp difference at the beginning of the
night.


------------------------------------------------------

Wednesday, January 19, 2011

Nightlog 2011-01-18

SA: Amanda
SO: Zolisa
Others: Siphelo, Patrick, Paul, Garith

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


No observations due to work on tracker.

-----------------------------------------

Tuesday, January 18, 2011

Nightlog 2010-01-17

SA: Amanda
SO: Zolisa
Others: Siphelo, Charl, Hamish, Jonathan, Green, Keith

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


. Mirror alignment work at night start. Tracker failures prevented us
from getting onsky and taking any data.

-----------------------------------------

Monday, January 17, 2011

Nightlog 2011-01-16

SA: Amanda
SO: Zolisa
Others: -

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Clear at night start, decent seeing. Closed for high humidity shortly
after midnight. Took some BCAM open-loop rho data for Tim.

-----------------------------------------

Sunday, January 16, 2011

Nightlog 2011-01-15

SA: Amanda
SO: Zolisa
Others: Jonathan

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


. Clear night with moderate seeing (1.5"-2"). Took one VCAM tile
during 1.15" seeing. Otherwise, took BCAM open-loop pointing data
for Tim and MASS/DIMM data all night.

-----------------------------------------

Saturday, January 15, 2011

Nightlog 2011-01-14

SA: Amanda
SO: Zolisa
Others: virtual Tim

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Took a small amount of BCAM and VCAM data in moderate seeing at night
start. Spent some time troubleshooting MASS/DIMM. High humidity and
ridge clouds claimed most of the night.

-----------------------------------------

Friday, January 14, 2011

Nightlog 2011-01-13

SA: Amanda
SO: Zolisa
Others: Jonathan, virtual Tim

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


No data taken. The humidity roller coaster allowed us to briefly take
some MASS/DIMM data and align mirrors in the middle of the night.
Otherwise, we were closed.

-----------------------------------------

Thursday, January 13, 2011

Nightlog 2011-01-12

SA: Amanda
SO: Zolisa
Others: Hamish, Jonathan

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Mostly clear night with poor seeing; night ended early due to tracker
failures. Took one set of VCAM data to allow plate scale calculation
and some BCAM open-loop rho pointing data.

-----------------------------------------

Wednesday, January 12, 2011

Nightlog 2011-01-11

SA: Paul
SO: Siphelo
Others: Zolisa, Amanda, Eben, Ockert, Chris, Martin, Johnathan,

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- We had a successful evening getting VCAM and BCAM parfocal. The VI stage was
moved back on dummy RSS. There was no Z adjustment needed for the VI stage. We
found that with VCAM in focus we could get BCAM focused using the FLI focus on
BCAM. Some further tests, which probably needs verification, showed that
the VCAM focus is fairly close to the theoretical track and not much
telescope focus offset is needed to get VCAM into focus.

- The weather conditions did not allow us to do more observations to
further check our initial findings.


------------------------------------------------------

Tuesday, January 11, 2011

Nightlog 2011-01-10

SA: Paul
SO: Siphelo
Others: Garith

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Seeing was not good enough for IQ. Did some open loop tracking observations.
We closed around midnight due to high humidity.

------------------------------------------------------

Monday, January 10, 2011

Nightlog 2011-01-09

SA: Paul
SO: Siphelo
Others: Garith, Ardhisha, Charl, Peter

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Opened up for alignment but it was cloudy all through the night and the
humidity forced us to close early.

------------------------------------------------------

Sunday, January 9, 2011

Nightlog 2011-01-08

SA: Paul
SO: Siphelo
Others: Garith, Ardhisha, Charl, Peter

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- The seeing was not good enough for IQ observations. We had cloud around at
the start of the evening, which returned just after 2.

- The tracker A-G motor was overheated when we got to the telescope. Charl and
Ardhisha came up to look at the alignment issue we had last night and also took
care of the overheating. The tracker behaved for the rest of the night.

- We had some alignment issues with the reference mirror not being found. Peter
came up to look at it.

------------------------------------------------------

Saturday, January 8, 2011

Nightlog 2011-01-07

SA: Paul
SO: Siphelo
Others: Garith, Ardhisha, Eben, Micheal

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Seeing conditions were not good enough for IQ observations. Continued with
the image drift observations.

- We closed early due a problem with the CCAS instrument when we tried a
re-alignment.

------------------------------------------------------

Friday, January 7, 2011

Nightlog 2011-01-06

SA: Paul
SO: Siphelo
Others: Garith, Hitesh

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- It was an eventful evening with good seeing and some tracker overheating
issues. We managed to get good seeing images with BCAM but struggled to get
good focus on VCAM.

- We had ~1'' seeing at one stage which we missed due the tracker overheating
problem, but we may have one or two images with ~1'' seeing on BCAM.

- At the end of the night the tracker x-lower motor overheated and we closed up
after that.

------------------------------------------------------

Thursday, January 6, 2011

Nightlog 2011-01-05

SA: Paul
SO: Siphelo
Others: Garith

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------

- Weather was good with low humidity for most of the night. Made some image
drift observations at a few azimuths.

- Had some alignment issues.

------------------------------------------------------

Wednesday, January 5, 2011

Nightlog 2011-01-04

SA: Petri
SO: Patrick
Others: Bruno, Paul, Garith, Vaisanen family, Charl

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


- Finished pointing model tests. Narrowed down the abnormal
pointings to Azimuth range of 357.5 to 360 only.

- Took two sets of open-loop rotation/tracking tests, before
overheating of tracker Rho motor stopped us before 3 am.

-----------------------------------------

Tuesday, January 4, 2011

Nightlog 2011-01-03

SA: Petri
SO: Patrick
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


- High humidity, never opened.

-----------------------------------------

Monday, January 3, 2011

Nightlog 2011-01-02

SA: Petri
SO: Patrick
Others: Bruno

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


- Night lost to humidity

-----------------------------------------

Sunday, January 2, 2011

Nightlog 2011-01-01

SA: Petri
SO: Patrick
Others: Bruno, Charl, Vaisanen family, Eben and Keith early on.

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


- Pointing model tests for most of the night. High humidity, but
steadily just below closing limit for the whole night.

-----------------------------------------

Saturday, January 1, 2011

Nightlog 2010-12-31

SA: Petri
SO: Patrick
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


- Wishing a Happy New Year 2011 for the SALT community!

And hoping it will be sunnier than the bleak weather conditions
here :) Lost night to rain.


-----------------------------------------